a tugboat goes 180 miles upstream in 10 hours. the return trip downstream takes 5 hours. find the speed of the tugboat without a current and the speed of the current.
Answer by scott8148(6628)

t - c = 180/10

t + c = 180/5

add the equations to eliminate c

solve for t

substitute back to find c